# Game-RL: Synthesizing Multimodal Verifiable Game Data to Boost VLMs’ General Reasoning

### v6

@@ 1 Introduction
Table 1: Our GameQA dataset extends RL training scenarios for VLMs to the domain of video games, providing diverse verifiable game tasks along with controllable difficulty. Comparison of existing visual reasoning datasets. Size means sum of the number of VQA pairs in train set and test set. Ration. Annot. means has annotated reasoning process. 3D Scene means some games of it are 3D scene. Detailed introduction can be found in Section 6.
@@ 5.4 Scaling effect of game diversity on generalization
With around 5,000 total training samples, we trained the Qwen2.5-VL-3B on GameQA subsets with 4, 10, and 20 distinct games (Table 7). The results (Figure 5) show a positive correlation between game diversity and generalization ability. This suggests that scaling game diversity makes better generalization, enabling the model to acquire more robust visual understanding and reasoning abilities.
@@ 5.5 Qualitative analysis
To confirm that GRPO substantially enhances visual perception and text reasoning abilities of models, we manually analyzed 790 cases randomly sampled from the results of InternVL2.5-8B, containing responses before and after GRPO. The results (Figure 5) confirm that after GRPO, the model demonstrates improved visual recognition of image elements and performs more precise reasoning. More statistics and cases are in Appendix 9.2 and 9.3.
@@ 8 Limitations and future work
(a) Number of distinct games
(b) Number of training task instances
Figure 7: The scaling effect of GameQA on unseen out-of-domain games
@@ 9.2 Error types analysis
9.3 More results and analysis on the impact of training game diversity and task instance diversity on model generalization
Game diversity While the model trained on 20 games achieved the highest performance gain on both unseen out-of-domain games and general visual benchmarks, interestingly, we observed that the model trained on 4 games (out-of-domain games +1.20%, general +0.66%) slightly outperformed the one trained on 10 games (out-of-domain games +0.96%, general +0.63%) on both benchmarks. We speculate that this non-linear improvement may result from cases where a smaller but well-chosen set of games provides a comparable amount of key generalization signals as a larger set, or that certain random factors affected the performance of specific game combinations.
Task instance diversity
Figure 11: The scaling effect of GameQA with an increasing number of task instances.
We further investigated how the task instance diversity influence model generalization using the Qwen2.5-VL-3B model. We compared three configurations: (a) 5k samples for 1 epoch, (b) 1k samples for 5 epochs, and (c) 0.5k samples for 10 epochs, ensuring consistent total training steps and batch sizes across all settings. On general visual benchmarks, the model trained with 5k samples for 1 epoch performed best, with a +1.2% improvement (Figure 11). This suggests that scaling game task instances diversity makes better generalization. Extended data scaling experiment is in Appendix 11.3.
On out-of-domain games in GameQA, unlike on general visual benchmarks, we observed a counterintuitive trend: as the number of training samples per epoch decreased and the number of training epochs increased, the model’s generalization performance improved, as shown in Figure 7. The best out-of-domain gain (+2.5%) was achieved by the model trained on 0.5k samples for 10 epochs, outperforming both the 1k $\times$ 5 epochs (+2.2%) and 5k $\times$ 1 epoch (+1.8%) configurations. This suggests that for near-domain tasks closely related to the training games, repeatedly training on a smaller but diverse dataset may help the model better capture task-specific patterns and nuances, leading to improved generalization.
@@ 10.4 Inference and evaluation
  • Evaluating the top three model checkpoints based on the validation set performance and averaging the evaluation results This is employed for the comparative training shown in Table 4 and the analytic experiments of Section 5.4, Appendix 9.3 and Appendix 11.2. This approach was chosen because the performance differences are often subtle and susceptible to randomness in the training process. Averaging across multiple top checkpoints helps smooth out training noise, enhances the robustness and sensitivity of the evaluation, and more accurately reflects the differences under various experimental settings. This method also reduces the impact of chance results, making our experimental conclusions more reliable and representative.
@@ 12.3 Time spent on the main steps of Code2Logic for each game
Figure 13 illustrates the estimated time spent on implementing the main steps of the Code2Logic approach across all the 30 games in the GameQA dataset. The time ranges from a minimum of 4 hours to a maximum of 12 hours, with an average of 7.5 hours per game.

### v8

@@ 1 Introduction
Table 1: Our GameQA dataset extends RL training scenarios for VLMs to the domain of video games, providing diverse verifiable game tasks along with controllable difficulty. Size means sum of the number of VQA pairs in train set and test set. Ration. Annot. means has annotated reasoning process. 3D Scene means some games of it are 3D scene. Detailed introduction can be found in Section 6.
@@ 2.4 Data quality assessment and data augmentation
Figure 3: Overview of the GameQA dataset. The 30 games in GameQA can be classified into four categories based on the core abilities required to solve game tasks. Appendix 13.2 provides definitions of the four game categories. Games chose as Out-of-Domain are not used for training; instead, they are used to test the generalization performance after the model has been trained on In-Domain games.
3 The GameQA dataset
@@ 5.4 Scaling effect of game diversity on generalization
With around 5,000 total training samples, we trained the Qwen2.5-VL-3B on GameQA subsets with 4 and 20 distinct games (Table 7). The results (Figure 6) show a positive correlation between game diversity and generalization ability. This suggests that scaling game diversity makes better generalization, enabling the model to acquire more robust visual understanding and reasoning abilities.
@@ 5.5 Qualitative analysis
To confirm that GRPO substantially enhances visual perception and text reasoning abilities of models, we manually analyzed 790 cases randomly sampled from the results of InternVL2.5-8B, containing responses before and after GRPO. The results (Figure 6) confirm that after GRPO, the model demonstrates improved visual recognition of image elements and performs more precise reasoning. More statistics and cases are in Appendix 9.2.
@@ 10.4 Inference and evaluation
  • Evaluating the top three model checkpoints based on the validation set performance and averaging the evaluation results This is employed for the comparative training shown in Table 4 and the analytic experiments of Section 5.4, Appendix LABEL:sec:more_results_on_diversity and Appendix 11.2. This approach was chosen because the performance differences are often subtle and susceptible to randomness in the training process. Averaging across multiple top checkpoints helps smooth out training noise, enhances the robustness and sensitivity of the evaluation, and more accurately reflects the differences under various experimental settings. This method also reduces the impact of chance results, making our experimental conclusions more reliable and representative.
@@ 12.3 Time spent on the main steps of Code2Logic for each game
Figure 11 illustrates the estimated time spent on implementing the main steps of the Code2Logic approach across all the 30 games in the GameQA dataset. The time ranges from a minimum of 4 hours to a maximum of 12 hours, with an average of 7.5 hours per game. Given that all annotators were undergraduates and that ten of them completed only a single game, the time required for a more experienced annotator could be significantly less than this average.